OpenStudy (anonymous):

Find the slope and y-intercept of the line 6x=5y+35

OpenStudy (anonymous):

slope=6/5 y-intercept=-7

jimthompson5910 (jim_thompson5910):

Solve for y and place the equation in y = mx+b form 6x=5y+35 6x-35=5y 5y=6x-35 y=6/5x-35/5 y=6/5x-7 Equation is now in y=mx+b form where m=6/5 and b=-7. So the slope is 6/5 and the y-intercept is (0,-7)